OpenStudy (anonymous):

The number of tickets sold each day for an upcoming performance of Handels Messiah is given by N(x)= -0.5x^2+12x+12 where x is the number of days since the concert was first announced. When will ticket sales peak and how many tickets will be sold that day!?

OpenStudy (anonymous):

you have to find the vertex of the parabola. this will be the maximum point at where it peaks. to find vertex, we first need to find axis of symmetry that is -b/(2a) so -12/(2(-0.5)) -12/-1 12 is the x value plug in 12 N(x)= -0.5x^2+12x+12 N(12)= -0.5(12)^2 +12(12) +12 N(12)= -0.5(144) +144+12 -72 +144+12 N(12)=84 the tickets will be at its peak on day 12, and they will sell 84 tickets

OpenStudy (anonymous):

Thank you!!!! :)

OpenStudy (anonymous):

no problem!